# MolFORM: Multi-modal Flow Matching for Structure-Based Drug Design

### v2

@@ 4.1 Experiment Setup
Our data processing strategy for DPO follows the methodology proposed in  Gu et al. (2024). We preprocess the dataset into a preference format $\mathcal{D}=\{(\mathbf{p},\mathbf{m}^{w},\mathbf{m}^{l})\}$, where $\mathbf{p}$ denotes the protein pocket, $\mathbf{m}^{w}$ the preferred ligand, and $\mathbf{m}^{l}$ the less preferred one. For each pocket, we sample two candidate ligands and assign preference based on a user-defined reward, mainly binding energy (e.g., Vina score). Since the affinity labels are continuous, we follow the strategy in Gu et al. (2024) and choose the molecule with the worst score as the dispreferred sample $\mathbf{m}^{l}$. This encourages a larger reward gap between $\mathbf{m}^{w}$ and $\mathbf{m}^{l}$, which has been shown to improve learning.
@@
MolFORM 0.4282 0.3419 0.3824 0.2898 0.3239 0.3583
Table 2: JSD Bond Length Comparisons across different methods.
Baselines.
We select all baseline models from CBGbench (Lin et al., 2024) for comparison with our method. Early structure-based drug design (SBDD) methods are built on voxel grids with deep neural networks, such as LiGAN (Ragoza et al., 2022), which generates atom voxelized density maps using variational autoencoders (VAE) and convolutional neural networks (CNNs), and 3DSBDD (Luo et al., 2022), which predicts atom types on grids with graph neural networks (GNNs) in an auto-regressive manner. The development of equivariant graph neural networks (EGNNs) enables direct generation of 3D atom positions, as seen in Pocket2Mol (Peng et al., 2022) and GraphBP (Liu et al., 2022a), which use auto-regressive strategies with normalizing flows. Diffusion-based methods such as TargetDiff (Guan et al., 2023), DiffBP (Lin et al., 2022), and DiffSBDD (Schneuing et al., 2024) generate atom types and positions using denoising diffusion probabilistic models. Recent methods incorporate domain knowledge to guide generation: FLAG (Zhang et al., 2023) and D3FG (Lin et al., 2023) use fragment motifs for coarse molecular generation; DecompDiff (Guan et al., 2024) uses scaffold and arm clustering with Gaussian process models for atom positions. More recent advances like MolCraft (Qu et al., 2024) and VoxBind (Pinheiro et al., 2024) apply new generative modeling strategies, including Bayesian flow networks and voxel-based diffusion with walk-jump sampling.
Figure 3: Comparing the distribution for distances of allatom (top row) and carbon-carbon pairs (bottom row) for reference molecules in the test set (gray) and model generated molecules (color).
4.2 Experiment result
Evaluation
We collect all generated molecules across 100 test proteins and evaluate generated ligands from three aspects: binding affinity, molecular properties and molecular structures. For target binding affinity and molecular properties, we present our results under the best setting as MolFORM. Following previous work, we utilize AutoDock Vina (Eberhardt et al., 2021) for binding affinity estimation. For binding affinity, we report the Vina Score, which evaluates the initially generated binding pose; the Vina Min, obtained after local energy minimization; and the Vina Dock, representing the lowest energy score from a global re-docking procedure using grid-based search. For molecular properties, we primarily report QED (Quantitative Estimate of Drug-likeness) and SA (Synthetic Accessibility). These results are summarized in Table 1. For molecular structures, we first evaluate several geometry-related properties following the setup in  Lin et al. (2024). These include (i) JSDBL and (ii) JSDBA, which measure the divergence in bond length and bond angle distributions between generated and reference molecules, reflecting structural realism. (iii) Ratiocca denotes the proportion of atoms with steric clashes—defined as van der Waals overlaps $\geq 0.4$Å—with protein atoms. (iv) Ratiocm captures the fraction of generated molecules that contain any such clashes. These results are also included in Table 1. (v) Root Mean Square Deviation (RMSD) is a standard metric to evaluate the structural deviation between two sets of atomic coordinates. Lower RMSD indicates higher structural consistency reflects the extent to which rigid fragments preserve their geometric integrity before and after force field optimization, which is shown in Figure 4. Besides, we report the ring portion for chemical assessment in Table 3. For substructure-level evaluation, we analyze the Jensen-Shannon divergence (JSD) between the bond length distributions of reference and generated molecules, as shown in Table 2. Here, we mainly compare with TargetDiff, which is highly similar to our proposed MolFORM.
Method 3 4 5 6 7 8
Ref. 0.0130 0.0020 0.2855 0.6894 0.0098 0.0003
LiGan 0.2238 0.0698 0.2599 0.4049 0.0171 0.0096
3DSBDD 0.2970 0.0007 0.1538 0.5114 0.0181 0.0116
GraphBP 0.0000 0.2429 0.1922 0.1765 0.1533 0.1113
Pocket2Mol 0.0000 0.1585 0.1822 0.4373 0.1410 0.0478
TargetDiff 0.0000 0.0188 0.2856 0.4918 0.1209 0.0298
DiffSBDD 0.2842 0.0330 0.2818 0.2854 0.0718 0.0193
DiffBP 0.0000 0.2195 0.2371 0.2215 0.1417 0.0707
FLAG 0.0000 0.0682 0.2716 0.5228 0.0996 0.0231
D3FG 0.0000 0.0201 0.2477 0.5966 0.0756 0.0283
DecompDiff 0.0302 0.0378 0.3407 0.4386 0.1137 0.0196
MolCraft 0.0000 0.0022 0.2494 0.6822 0.0489 0.0072
VoxBind 0.0000 0.0062 0.2042 0.7566 0.0232 0.0021
MolFORM 0.0000 0.0291 0.2846 0.4211 0.2033 0.0455
Table 3: Distribution of different ring sizes across various methods.
Result analysis
The visualization comparison result is shown in Figure 2. As shown in Table 1, we observed that MolFORM achieved comparable performance to our baseline model TargetDiff in binding affinity and molecular properties. Compared to Targetdiff, the model shows better performance in terms of Clash and Vina Dock. After further DPO fine-tuning, our model achieves state-of-the-art performance across nearly all metrics.
Although we use only binding affinity as the reward, DPO aligns the generation distribution toward higher-quality molecules while preserving the chemical priors from the pretrained model, thus avoiding overfitting to a single objective. As a result, metrics such as QED and SA can also improve. This is further supported by the enhanced quality of generated molecular substructures in Figure 3 and Figure 4. Additionally, we observe that the diversity of generated molecules remains at 0.77 after DPO fine-tuning, indicating that our approach successfully avoids mode collapse while improving generation quality.
Metric TargetDiff TargetDiff-DPO MolFORM MolFORM-DPO
Vina Score ($\downarrow$) -5.47 -5.58 -5.42 -6.16
Vina Min ($\downarrow$) -6.39 -6.59 -6.42 -7.18
QED ($\uparrow$) 0.46 0.47 0.48 0.50
SA ($\uparrow$) 0.60 0.61 0.60 0.65
Table 4: Comparison of model performance before and after DPO fine-tuning. The reported numbers are average values here. The results of vanilla TargetDiff are obtained by retraining the model.
@@ 4.2 Experiment result
We conducted ablation studies on the Chamfer DPO loss component in Table 5. Our experiments demonstrate that incorporating the Chamfer distance into the DPO objective significantly enhances the model’s ability to preserve molecular geometric fidelity.
Metric Chamferw Chamferwo
@@
Table 5: Ablation study: Comparison of model performance with (Chamferw) and without (Chamferwo) Chamfer loss in Vanilla Multi-model flow matching.
Sampling efficiency
We achieve SOTA sampling performance. While it takes on average 3428s and 6189s for TargetDiff and DecompDiff to generate 100 samples respectively, our model only uses 69s. Thanks to the efficiency of flow matching, our sampling requires only 100 steps, compared to the 1000 steps typically used in diffusion models.
Training details
The model is trained via the gradient descent method Adam Kingma and Ba (Kingma & Ba, 2014) with init learning $\text{rate}=5\times 10^{-5}$, betas=(0.95, 0.999), batch size=16, and clip gradient norm=8. During training, we apply data augmentation by adding small Gaussian noise (standard deviation of 0.1) to the protein atom coordinates. Additionally, we use an exponential learning rate decay with a decay factor of 0.6 and a minimum threshold set at $1\times 10^{-8}$. For the DPO model training, a smaller learning rate was applied as $\text{rate}=5\times 10^{-8}$ with a decay factor of 0.6 and a minimum threshold set at $1\times 10^{-11}$ and the beta dpo parameter was set as 5.
We trained our model on two NVIDIA GH200 GPUs. The base model training converged within 12 hours and 100k steps. Moreover, the DPO model training utilized 4 hours and 20k steps.
Figure 4: Median RMSD for rigid fragment before and after the force-field optimization.
5 Future Work
Our model builds upon  Guan et al. (2023) and achieves comparable or even superior performance across multiple evaluation metrics, indicating that the proposed MultiFlow-based modeling is a promising direction for further exploration. Beyond DPO, other fine-tuning strategies and multi-objective optimization techniques also hold promise for further advancing the performance of our current model. Moreover, we employ uniform corruption for discrete flow matching, as masked variants unexpectedly underperformed in our molecular generation task—contrasting with their success in other scientific domains. Similarly, SDE sampling via Euler-Maruyama yielded inferior results compared to ODE integration, though the specific hyperparameters for stochastic sampling may require further tuning. These findings highlight the importance of tailoring both corruption strategies and sampling algorithms to the specific requirements of SBDD task.
6 Conclusions
In this work, we propose to use a novel Multi flow based generation framework MolFORM for the protein-specific molecule generation. Extensive experiments on the CrossDocked2020 benchmark demonstrate the strong performance of MolFORM. Our framework can be further combined with the Direct Preference Optimization (DPO) process, achieving state-of-the-art performance across almost all evaluation metrics. Measurements on molecular substructures indicate that the improvement is not merely due to overfitting on specific metrics, but reflects a holistic enhancement of generation quality. Moreover, our comparison with TargetDiff demonstrates that this DPO-based approach offers greater potential than unimodal fine-tuning.
Acknowledgements
This work is supported by the NSF 2340011. We would like to thank the reviewers for taking the time and effort necessary to review the manuscript. We sincerely appreciate Dr. Chengyue Gong for valuable comments.
Impact Statement
We introduce a multi-flow DPO framework that enables preference-guided generation over both discrete atom types and continuous molecular coordinates. Beyond SBDD, the generality of our framework suggests potential extensions to broader domains of computer-aided design, including small molecule generation, material discovery, and chip layout optimization. We stress the importance of ensuring the responsible deployment of this technology and caution against its misuse in harmful or unethical contexts.

### v3

@@ 4.1. Experiment Setup
Our data processing strategy for DPO follows the methodology proposed in  Gu et al. (2024). We preprocess the dataset into a preference format $\mathcal{D}=\{(\mathbf{p},\mathbf{m}^{w},\mathbf{m}^{l})\}$, where $\mathbf{p}$ denotes the protein pocket, $\mathbf{m}^{w}$ the preferred ligand, and $\mathbf{m}^{l}$ the less preferred one. For each pocket, we sample two candidate ligands and assign preference based on a user-defined reward, mainly binding energy (e.g., Vina score). Since the affinity labels are continuous, we follow the strategy in Gu et al. (2024) and choose the molecule with the worst score as the dispreferred sample $\mathbf{m}^{l}$, which encourages a larger reward gap between $\mathbf{m}^{w}$ and $\mathbf{m}^{l}$.
RL dataset
Distinct from static offline datasets, we adopt an on-policy data collection strategy. Using the processed CrossDocked2020 pockets as the environment, we dynamically generate candidate ligands via policy rollouts during each training iteration. Valid molecules are evaluated using specified reward functions (e.g., Vina, QED, SA). To stabilize training, rewards are normalized within each pocket—centered, scaled, clipped to [1,1][-1,1], and mapped to [0,1][0,1]—forming temporary tuples used exclusively for the current parameter update.
RL rollout and scoring.
We maintain an EMA anchor policy (decay $0.995$) for sampling and as the reference network. For each update, we sample B=4B{=}4 pockets and generate K=8K{=}8 ligand candidates per pocket using a 100-step log-time multi-flow sampler (atom-count prior; discrete temperature $0.01$; discrete noise $1.0$). We evaluate each generated ligand by AutoDock Vina and compute a normalized synthetic accessibility (SA) score. Given tuples $(p,m_{p,i},r_{p,i})$, we sample $t\sim\mathcal{U}(0,1)$ and apply the same forward corruptions as pretraining. We form implicit positive/negative branches by mixing current and anchor predictions with $\beta=\beta_{\mathrm{discrete}}=0.3$, and minimize the reward-weighted DiffusionNFT objective using Adam (lr $5\times 10^{-7}$) with gradient clipping $8.0$.
@@
Baselines.
We select several baseline models from CBGbench (Lin et al., 2024) for comparison with our method. Early structure-based drug design (SBDD) methods are built on voxel grids with deep neural networks, such as LiGAN (Ragoza et al., 2022), which generates atom voxelized density maps using variational autoencoders (VAE) and convolutional neural networks (CNNs), and 3DSBDD (Luo et al., 2022), which predicts atom types on grids with graph neural networks (GNNs) in an auto-regressive manner. The development of equivariant graph neural networks (EGNNs) enables direct generation of 3D atom positions, as seen in Pocket2Mol (Peng et al., 2022) and GraphBP (Liu et al., 2022a), which use auto-regressive strategies with normalizing flows. Diffusion-based methods such as TargetDiff (Guan et al., 2023), DiffBP (Lin et al., 2022), and DiffSBDD (Schneuing et al., 2024) generate atom types and positions using denoising diffusion probabilistic models. Recent methods incorporate domain knowledge to guide generation: FLAG (Zhang et al., 2023) and D3FG (Lin et al., 2023) use fragment motifs for coarse molecular generation; DecompDiff (Guan et al., 2024) uses scaffold and arm clustering with Gaussian process models for atom positions. More recent advances like MolCraft (Qu et al., 2024) and VoxBind (Pinheiro et al., 2024) apply new generative modeling strategies, including Bayesian flow networks and voxel-based diffusion with walk-jump sampling.
We additionally include three recent strong baselines: TAGMol (Dorna et al., 2024), which uses target-aware gradient guidance to steer diffusion sampling toward better affinity and auxiliary properties; DecompOpt (Zhou et al., 2024a), a controllable decomposed diffusion framework for structure-based molecular optimization; and MolJO (Qiu et al., 2024), a gradient-guided Bayesian-update approach that jointly optimizes discrete atom types and continuous coordinates.
Figure 3. Comparing the distribution for distances of allatom (top row) and carbon-carbon pairs (bottom row) for reference molecules in the test set (gray) and model generated molecules (color).
Figure 4. Visualizations of reference molecules and generated ligands for protein pockets (4yhj) generated by Reference, Targetdiff, MolFORM, MolFORM-DPO and MolFORM-RL. Vina score, QED, and SA are reported below.
4.2. Experiment result
Evaluation
We collect all generated molecules across 100 test proteins and evaluate generated ligands from three aspects: binding affinity, molecular properties and molecular structures. For target binding affinity and molecular properties, we present our results under the best setting as MolFORM. Following previous work, we utilize AutoDock Vina (Eberhardt et al., 2021) for binding affinity estimation. For binding affinity, we report the Vina Score, which evaluates the initially generated binding pose; the Vina Min, obtained after local energy minimization; and the Vina Dock, representing the lowest energy score from a global re-docking procedure using grid-based search. For molecular properties, we primarily report QED (Quantitative Estimate of Drug-likeness) and SA (Synthetic Accessibility). These results are summarized in Table 1. For molecular structures, we first evaluate several geometry-related properties following the setup in  Lin et al. (2024). These include (i) JSDBL and (ii) JSDBA, which measure the divergence in bond length and bond angle distributions between generated and reference molecules, reflecting structural realism. (iii) Ratiocca denotes the proportion of atoms with steric clashes—defined as van der Waals overlaps $\geq 0.4$Å—with protein atoms. (iv) Ratiocm captures the fraction of generated molecules that contain any such clashes. These results are also included in Table 1. Further structural evaluations, including Root Mean Square Deviation (RMSD), ring size distributions, and substructure-level bond length JSD, are detailed in the Appendix (Figure 7, Table 3, and Table 4).
Figure 5. Effectiveness of the Confidence Head. Selecting a smaller ratio of top-confidence samples yields significantly better (lower) Vina scores.
Figure 6. Median Vina energy for different generated molecules (TargetDiff vs. MolCraft vs. MolJO vs. MolFORM RL) across 100 testing binding targets. Binding targets are sorted by the median Vina energy of generated molecules. Lower Vina energy means a higher estimated binding affinity.
Result analysis
As summarized in Table 1, our base model MolFORM demonstrates strong generative capabilities. While achieving comparable binding affinity to the diffusion baseline TargetDiff, MolFORM significantly outperforms it in structural validity and diversity. Specifically, MolFORM reduces the steric clash ratio (Ratiocca) from 0.0483 to 0.0310 and achieves a higher diversity score (0.78 vs. 0.72), indicating its ability to explore a broader chemical space with physically plausible structures.The introduction of preference alignment yields substantial improvements. MolFORM-DPO excels in balancing generation quality and diversity. It achieves a remarkable reduction in steric clashes (0.0188) and improves the Vina Score to -6.16, while maintaining a high diversity of 0.77.
Furthermore, MolFORM-RL achieves state-of-the-art performance in binding affinity, with a Vina Score of -7.60, surpassing recent strong baselines such as TAGMol (-7.02) and MolJO (-7.52). This suggests that our alignment strategies successfully steer the generative flow toward chemically favorable manifolds, preserving the intrinsic chemical priors learned during pre-training. Meanwhile, MolFORM-RL maintains a high diversity score of 0.75, highlighting a key distinction from other fine-tuning approaches that often improve affinity at the cost of diversity.
Finally, we validate the effectiveness of the auxiliary Confidence Head. As illustrated in Figure 5, the predicted confidence scores exhibit a strong correlation with the ground-truth structural quality (measured by RMSD). This indicates that the confidence head serves as a reliable estimator, enabling efficient ranking and filtration of high-fidelity candidates.
4.3. Ablation study on Reward design
We employed three distinct reward formulations. The first one is formulated as a linear combination of normalized Vina Score and SA as follows:
[…]
where $\text{Clip}(\cdot,\cdot,\cdot)$ denotes the clipping operation. The second type combines QED and Vina score, while the third corresponds to QED and SA. We observed that the first two types yield sustained performance gains; however, the third type does not lead to significant improvements in the Vina score. Detailed reward curves illustrating the online RL training process are provided in the Figure 8.
Table 2. Ablation study on different reward formulations. Reward 1 (Vina + SA) achieves the best binding affinity, while Reward 2 improves QED. Reward 3 shows limited improvement on Vina score.
Metric Reward 1 Reward 2 Reward 3
(Vina + SA) (QED + Vina) (QED + SA)
Vina Score ($\downarrow$) -7.60 -7.08 -4.38
Vina Min ($\downarrow$) -8.37 -8.13 -5.84
QED ($\uparrow$) 0.50 0.60 0.56
SA ($\uparrow$) 0.68 0.56 0.66
5. Conclusions
In this work, we introduce MolFORM, a multimodal flow matching framework for protein-specific molecular generation that jointly models discrete atom types and continuous 3D coordinates. We further show that online reinforcement learning provides a powerful mechanism for aligning flow-based generative models with biochemical objectives. On the CrossDocked2020 benchmark, MolFORM-RL achieves state-of-the-art binding affinity. Moreover, our reinforcement learning framework holds strong potential for extension to other structure-based drug design (SBDD) generative models.
Appendix A Algorithm
Algorithm 1 Forward-Process Negative-aware Online Fine-tuning for Conditional Diffusion-based Ligand Generation
1:Pretrained conditional diffusion/flow policy $v_{\mathrm{ref}}$; pocket dataset $\mathcal{C}$ (conditions $c$); raw reward function $r_{\mathrm{raw}}(\mathbf{x}_{0},c)\in\mathbb{R}$; rollout size $K$; mixing coefficient $\beta_{\mathrm{NFT}}$; learning rate $\lambda$; EMA update ratio $\eta_{i}$; diffusion schedule / forward corruption operator $q_{t|0}$.
2:Notation: a ligand sample is $\mathbf{x}_{0}=(\mathbf{s}_{0},\mathbf{r}_{0})$ (atom types $\mathbf{s}_{0}$ and 3D coordinates $\mathbf{r}_{0}$).
3:Initialize: sampling policy $v_{\mathrm{old}}\leftarrow v_{\mathrm{ref}}$, training policy $v_{\theta}\leftarrow v_{\mathrm{ref}}$, replay buffer $\mathcal{D}\leftarrow\emptyset$.
4:for iteration $i=1,2,\dots$ do
5:  for each sampled pocket condition $c\sim\mathcal{C}$ do $\triangleright$ Rollout / Data Collection
6:   Sample $K$ ligand structures $\{\mathbf{x}_{0}^{k}\}_{k=1}^{K}\sim\pi_{\mathrm{old}}(\cdot\,|\,c)$ using any black-box solver.
7:   Compute rewards $r_{\mathrm{raw}}^{k}\leftarrow r_{\mathrm{raw}}(\mathbf{x}_{0}^{k},c)$ for $k=1..K$.
8:   Group-normalize rewards: $r_{\mathrm{norm}}^{k}\leftarrow r_{\mathrm{raw}}^{k}-\frac{1}{K}\sum_{j=1}^{K}r_{\mathrm{raw}}^{j}$.
9:   Convert to optimality probability $r^{k}\in[0,1]$:
$r^{k}\leftarrow\frac{1}{2}+\frac{1}{2}\,\mathrm{clip}\!\left(\frac{r_{\mathrm{norm}}^{k}}{Z_{c}},-1,1\right),$
where $Z_{c}>0$ is a normalizer (e.g., global reward std).
10:   Add tuples to buffer: $\mathcal{D}\leftarrow\mathcal{D}\cup\{(c,\mathbf{x}_{0}^{k},r^{k})\}_{k=1}^{K}$.
11:  end for
12:  for each minibatch $\{(c,\mathbf{x}_{0},r)\}\subset\mathcal{D}$ do $\triangleright$ Policy Optimization on Forward Process
13:   Sample time $t\sim\mathcal{U}(0,1)$ and noises (Gaussian for coordinates; discrete corruption noise for atom types).
14:   Forward corruption: sample noisy state $\mathbf{x}_{t}\sim q_{t|0}(\cdot\,|\,\mathbf{x}_{0})$.
15:   Compute the standard diffusion/flow regression target $\mathbf{v}$ under your parameterization (e.g., velocity/flow/score target).
16:   Implicit positive branch:
$v_{\theta}^{+}(\mathbf{x}_{t},c,t)\leftarrow(1-\beta_{\mathrm{NFT}})\,v_{\mathrm{old}}(\mathbf{x}_{t},c,t)+\beta_{\mathrm{NFT}}\,v_{\theta}(\mathbf{x}_{t},c,t).$
17:   Implicit negative branch:
$v_{\theta}^{-}(\mathbf{x}_{t},c,t)\leftarrow(1+\beta_{\mathrm{NFT}})\,v_{\mathrm{old}}(\mathbf{x}_{t},c,t)-\beta_{\mathrm{NFT}}\,v_{\theta}(\mathbf{x}_{t},c,t).$
18:   Update parameters by minimizing the negative-aware forward loss:
$\theta\leftarrow\theta-\lambda\nabla_{\theta}\Big(r\|v_{\theta}^{+}-\mathbf{v}\|_{2}^{2}+(1-r)\|v_{\theta}^{-}-\mathbf{v}\|_{2}^{2}\Big).$
19:  end for
20:  $\triangleright$ Online EMA update of sampling policy (off-policy)
21:  $\theta_{\mathrm{old}}\leftarrow\eta_{i}\,\theta_{\mathrm{old}}+(1-\eta_{i})\,\theta$;  $\mathcal{D}\leftarrow\emptyset$.
22:end for
23:Output: fine-tuned conditional policy $v_{\theta}$.
Appendix B Additional Experimental Results
In this section, we provide detailed structural and chemical assessments of the generated molecules to validate their geometric integrity and distribution consistency.
Structural Deviation Analysis
Figure 7 illustrates the Median Root Mean Square Deviation (RMSD) for rigid fragments before and after force-field optimization. RMSD serves as a standard metric to evaluate structural deviation; lower values indicate higher structural consistency. The results reflect the extent to which rigid fragments preserve their geometric integrity, demonstrating that our generated structures are chemically stable and require minimal adjustment during optimization.
Chemical Property Distributions
We further evaluate the chemical validity of the generated molecules through bond length distributions and ring size analysis. Table 3 presents the Jensen-Shannon Divergence (JSD) of bond lengths for various bond types compared to the reference dataset. Lower JSD scores indicate that the generated bond lengths closely match the ground truth distribution. Additionally, ring size distributions are detailed in Table 4, confirming that the model generates chemically plausible ring structures.
Figure 7. Median RMSD for rigid fragment before and after the force-field optimization.
Table 3. JSD Bond Length Comparisons across different methods.
@@ Appendix B Additional Experimental Results
MolFORM 0.4852 0.3876 0.3830 0.3919 0.5496 0.4372
Table 4. Proportion (%) of different ring sizes in reference and generated ring-structured molecules, where 3-Ring denotes three-membered rings and the like.
3-Ring 4-Ring 5-Ring 6-Ring
Reference 4.0 0.0 49.0 84.0
Train 3.8 0.6 56.1 90.9
AR 50.8 0.8 35.8 71.9
Pocket2Mol 0.3 0.1 38.0 88.6
FLAG 3.1 0.0 39.9 84.7
TargetDiff 0.0 7.3 57.0 76.1
DecompDiff 9.0 11.4 64.0 83.3
IPDiff 0.0 6.4 51.0 83.7
MolCRAFT 0.0 0.6 47.0 85.1
DecompOpt 6.8 11.8 61.4 89.8
TAGMol 0.0 8.5 62.5 82.6
MolJO 0.0 0.3 44.4 97.6
MolFORM 0.0 3.2 52.0 84.4
Ring size
Furthermore, we analyze the ring size distribution to assess the topological realism of the generated molecules. Table 4 compares the frequencies of different ring sizes (e.g., 3-, 4-, 5-, and 6-membered rings) against the reference dataset. The results demonstrate that our model generates chemically plausible ring structures, avoiding the over-generation of unstable small rings or unrealistic large fused systems often observed in baseline methods.
@@
We conducted ablation studies on the Chamfer DPO loss component in Table 5. Our experiments demonstrate that incorporating the Chamfer distance significantly enhances the model’s ability to preserve molecular geometric fidelity.
Table 5. Ablation study: Comparison of model performance with (Chamfer-w) and without (Chamfer-wo) Chamfer loss in Vanilla Multi-model flow matching.
Metric Chamfer-w Chamfer-wo
@@
Abalation on DPO training
Table 6. Comparison of model performance before and after DPO fine-tuning. The reported numbers are average values here. The results of vanilla TargetDiff are obtained by retraining the model.
Metric TargetDiff TargetDiff-DPO MolFORM MolFORM-DPO
Vina Score ($\downarrow$) -5.47 -5.58 -5.42 -6.16
Vina Min ($\downarrow$) -6.39 -6.59 -6.42 -7.18
QED ($\uparrow$) 0.46 0.47 0.48 0.50
SA ($\uparrow$) 0.60 0.61 0.60 0.65
Training detail
To ensure training stability during the alignment phase, we employed significantly lower learning rates compared to the pre-training stage. Specifically, the learning rate was set to $5\times 10^{-8}$ for DPO and $5\times 10^{-6}$ for online RL, which are considerably smaller than the base model’s learning rate of $5\times 10^{-5}$. For the MolFORM-RL configuration, we set the KL regularization coefficients $\beta$ and $\beta_{\text{discrete}}$ to 0.3 to prevent excessive deviation from the prior distribution. Additionally, we utilized an Exponential Moving Average (EMA) with a decay rate of 0.995 for target network updates and applied gradient clipping with a maximum norm of 8.0 to further stabilize the optimization process. Detailed reward curves illustrating the training stability are provided in the Appendix. We observed that employing a composite reward of QED, SA, and Vina score yields superior performance, effectively enhancing binding affinity while maintaining high drug-likeness.
Figure 8. Online RL reward curves for Reward 1 and Reward 2.